OpenStudy (anonymous):

Please Help Tyler1997 4.One of your friends sends you an email asking you to explain how all of the following expressions have the same answer. 3 sqwrt x^3 x^1/3*x^1/3*x^1/3 1/x^-1 11sqwrt x^5*x^4*x^2 Compose an email back assisting your friend and highlight the names of the properties of exponents when you use them.

OpenStudy (anonymous):

by 3sqwrtA you mean \[\sqrt[3]{A}\] ??? Cause that is called cubical root. well if so then cbrt(x^3) is a cubical root of the third power of x so it is equal to x. Samewise the second one, since x^(1/3)=cbrt(3) x(-1)=1/x so if you divide 1/(1/x) Then it is equal x x^a*x^b=x^(a+b) thus x^5*x^4*x^2=(x^5*x^4)*x^2=x^(5+4)*x^2=x^9*x^2=x^(9+2) Now the root of eleventh degree from x^11 is also equal to x. All those are equal to x so they are all equal to each other.